OpenStudy (anonymous):

last time I'll ask, it's hard so DON'T click if you know you can't solve it! H(x)=(x^2)+1 and K(x)=-(x^2)+4. If K(H)=0, what are the roots/solutions? a. ±i√3 b. ±i c. ±2 d. ±1 e. ±1, ±i√3

OpenStudy (cmtboy2):

d. ±1 i think

OpenStudy (anonymous):

please, please explain why..

OpenStudy (anonymous):

i think it's e but i'm not so sure

OpenStudy (cmtboy2):

can it be to answer

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

no, its only 1

OpenStudy (cmtboy2):

@eskawaiidesu can it be two Answer or no

OpenStudy (anonymous):

no, it can only be one

OpenStudy (cmtboy2):

ok

OpenStudy (cmtboy2):

um.... i wold go with C. But D. is also is the right answer to

OpenStudy (anonymous):

okay, i'll see..

OpenStudy (cmtboy2):

ok

OpenStudy (cmtboy2):

go with c

OpenStudy (cmtboy2):

what grad are you in

OpenStudy (anonymous):

I'm in college :/

OpenStudy (cmtboy2):

oh :)

OpenStudy (cmtboy2):

i have to go bye

OpenStudy (trojanpoem):

H(x)=(x^2)+1 and K(x)=-(x^2)+4 substitute H(x) instead of x to get K(H) K(H) = -(x^2+1)^2 + 4 K(H) = - (x^4 + 1 + 2x^2) + 4 K(H) = -x^4 - 1 - 2x^2 +4 = -x^4 -2x^2 + 3 Let K(H) = 0 X^4 + 2x^2 - 3 = 0 X^2 = 1 , x^2 = -3 Get the roots x = + or - 1 x = + or - sqrt(3)i I assume, you can use the quadratic formula -b +- sqrt(b^2 - 4ac)/2a Can't find the hard part you 're talking about..

OpenStudy (anonymous):

thank you

OpenStudy (trojanpoem):

Any time.